OpenStudy (ttop0816):

is this undefined?? sec(arcsec1/2) undefined -1/2 1/2 2

OpenStudy (ttop0816):

@superdavesuper could you help me with this question also??(:

OpenStudy (anonymous):

it is the sec of the angle whose sec is 0.5

OpenStudy (ttop0816):

@douglaswinslowcooper which would be 1/2 correct?

OpenStudy (anonymous):

Just so.

OpenStudy (superdavesuper):

but then since sec=1/cos; which means cos=2....so it is undefined :)

OpenStudy (ttop0816):

@superdavesuper right?? so it is undefined!

OpenStudy (superdavesuper):

well @douglaswinslowcooper eqn is right as well....but....im sticking w my ans though lol

OpenStudy (anonymous):

Good point, cannot have sec = 1/2 as cannot have cos = 2. My error.

OpenStudy (mathmale):

Guys, remember that the secant and inverse secant functions "undo" each other. Thus, sec(arcsec 1/2)=1/2.

OpenStudy (superdavesuper):

@mathmale but u evaluate the expression from the innermost bracket first, so i think it would become undefined BEFORE it got "undo" by the outer sec function. Just my opinion thou...

OpenStudy (ttop0816):

okay! sorry for asking but would it be 30 degrees if i simplify arccot sq.root3/3?

OpenStudy (superdavesuper):

@ttop0816 u may want to ask 4 more opinions on ur original prob. as for arccot(sqrt(3)/3)... let it be x, so cotx=sqrt3/3, so tanx=3/sqrt3 x=60

OpenStudy (loser66):

I am with mathmale, sec ( arcsec (something) ) = something.

OpenStudy (anonymous):

Don't tell me I was wrong about my being wrong!

OpenStudy (ttop0816):

ohhhh then it is 1/2?

OpenStudy (mathmale):

Yes.

OpenStudy (superdavesuper):

alright i will do a shout-out... @satellite73 @Hero @ganeshie8 @RadEn should the expression sec(arcsec1/2) be undefined or 1/2? thanx for ur hlep! :)

OpenStudy (loser66):

Let me help @ybarrap @whpalmer4 @wolfe8

OpenStudy (anonymous):

I was abused on a Mathematics section when someone came and asked for stories about abuse....

OpenStudy (superdavesuper):

good one @douglaswinslowcooper

OpenStudy (superdavesuper):

@GymnasticsGirl9052 there is no need to post ur question within someone else's. plz go post it as ur own. thank you :)

OpenStudy (anonymous):

Just kidding around, GG.

OpenStudy (mathmale):

Would someone please look up the graph of the secant function? If so, it'll become evident that the secant is always >= 1 or <=-1. Therefore, the correct answer to the problem posted by @ttop0816 is "undefined." My initial response (1/2) was wrong, and I apologize.

OpenStudy (loser66):

I don't agree!! they ask for the sec (arcsec (1/2) ) not sec (1/2) If you say so, so wolfram's answer is wrong???

OpenStudy (anonymous):

f[undefined] = undefined, I think.

OpenStudy (mathmale):

Can anyone come up with an angle whose secant is 1/2? Note that this question is equivalent to asking: Can anyone come up with an angle whose cosine is 2/1?

OpenStudy (raden):

hmmm... sec(arcsec(1/2)) let arcsec(1/2) is theta. if we draw it, looks the hypotenuse smallest then the leg of triangle. while in right triangle the hypotenuse always be longer, right ?

OpenStudy (mathmale):

RadEn...please come to a conclusion and share it with the rest of us.

OpenStudy (superdavesuper):

wish OS has a way to post a poll - it would be perfect for a question like this :)

OpenStudy (raden):

i want represented it in a diagram, but tool draw is not working now. to me, it is not exist. haha :P

OpenStudy (mathmale):

Or, paraphrased, there is no angle whose secant is (1/2). Would someone care to evaluate arcsec 2?

OpenStudy (superdavesuper):

one more round n i will call it quit on this... @jim_thompson5910 @agent0smith should the expression sec(arcsec1/2) be undefined or 1/2? thanx for ur hlep! :)

OpenStudy (ranga):

Undefined. arcsec(1/2) does not exist.

OpenStudy (agent0smith):

Yeah, the domain of arcsec is as stated above, x<-1 or x>1

OpenStudy (superdavesuper):

thank you @ranga! kind of disappointed that wolfram didnt check for bound in its calculation.

OpenStudy (superdavesuper):

thank you @agent0smith

OpenStudy (superdavesuper):

Thank you all....sorry i already gave out the medal on this earlier but u have my eternal thanks! :)

OpenStudy (agent0smith):

You can only "cancel out" the sec(arcsec ...) IF the ... is in the domain of the inner function. More simple example: cos(arccos 2) cannot equal 2, right? cos(...) can't be >1

OpenStudy (agent0smith):

But even still... http://www.wolframalpha.com/input/?i=cos%28arccos2%29

OpenStudy (superdavesuper):

:( wonder if we can report that to wolfram! lol

OpenStudy (agent0smith):

^i think it's because they do it over complex numbers, so... http://www.wolframalpha.com/input/?i=arccos2

OpenStudy (superdavesuper):

damn....i must have slept thro THAT part of my trig class!!! O^O thanks again @agent0smith

OpenStudy (agent0smith):

^no, i've never learned complex solutions to inverse trig functions either

OpenStudy (whpalmer4):

Yes, that's correct — ArcSec of a complex number... See http://mathworld.wolfram.com/InverseSecant.html

OpenStudy (whpalmer4):

I'm happy to say that I've avoided needing to know that stuff for more than 50 years, and hope to keep it that way for another 50 :-)

OpenStudy (whpalmer4):

but yes, a very interesting thing to figure out!

OpenStudy (superdavesuper):

hmmm.....thanks i guess @whpalmer4 lol so based on the wolfram description: if "Arcsecz being used for the multivalued function" then i guess the answer would be 1/2; but it is referring to the principal value only, then it would be undefined.... OK this is it.....i guess everyone is right then :)

OpenStudy (agent0smith):

@superdavesuper yep and it's very likely this is from a standard precalc/trig class, one that would not deal with complex number solutions

OpenStudy (superdavesuper):

we were more arguing on principle than anything else as we dont expect the "textbook" solution would come anywhere close to what are being discussed here :)

OpenStudy (loser66):

@superdavesuper 100% agree!! I was taught that a *b =0 --> a =0 or b =0 now I am taught that a*b =0 and no need to have either a =0 or b =0 (in other word a !=0 and b!=0)

OpenStudy (superdavesuper):

hi5 @Loser66 :) OK everybody goes home happy! lol im out of here - again thank you all for ur contribution!!

OpenStudy (whpalmer4):

let's just hope that no one had the misfortune of answering -1/2 or 2, no amount of "math lawyering" is going to get you a correct mark there :-)